what is the solution of 99th percentile?


1
Expert's answer
2021-06-07T17:27:25-0400



When we say that an individual’s test score was at the 99th percentile of the population, we mean that 99% of all population scores were below that score and 1% were above. 

Let "p" be a number between 0 and 1. The "(100p)"th percentile of the distribution of a continuous rv "X," denoted by "\eta(p)", is defined by


"p=F(\eta(p))=\displaystyle\int_{-\infin}^{\eta(p)}f(y)dy"

Acoording the expression "\eta(0.99)," the 99th percentile, is that value on the measurement axis such that "100(0.99)\%" of the area under the graph of "f(x)" lies to the left of "\eta(p)" and "100(1-0.99)\%" lies to the right.


For a example for the normal distribution the formula below is used to compute percentiles of a normal distribution.


"X=\mu+Z\sigma"

where "\mu" is the mean and "\sigma" is the standard deviation of the variable "X,"and "Z" is the value from the standard normal distribution for the desired percentile.

For the 99th percentile "\alpha=0.01" and "z_{\alpha}=2.3263."